the degree three polynomial f(x) with real coefficients and leading coefficient 1, has -3 and +4i among its roots. Express f(x) as the productof of linear and quadriatic polynomials with real coefficients
Answer by Edwin McCravy(20086)

the degree three polynomial f(x) with real coefficients and leading coefficient 1, has -3 and +4i among its roots. Express f(x) as the productof of linear and quadriatic polynomials with real coefficients

If a polynomial has all real coefficients and has A+Bi as a root,
then its conjugate A-Bi is also a root of the the polynomial.

So since the polynomial we want has +4i as a root, then it also
has -4i as a root. 

That's because +4i is really 0+4i and so its conjugate is 0-4i,
which is just -4i.

So we work the problem backwards:

We are to end up with:

    x = -3;        x = 4i         x = -4i

Before that we must have had this:

x + 3 = 0;    x - 4i = 0;    x + 4i = 0

Before that we must have had this:

            (x + 3)(x - 4i)(x + 4i) = 0

Multiplying the second and third parentheses together
using FOIL:

     (x + 3)(x² + 4ix - 4ix - 16i²) = 0

Canceling the +4ix and the -4ix

                 (x + 3)(x² - 16i²) = 0

Replacing i² by -1

                (x + 3)[x² -16(-1)] = 0  

Simplifying:

                   (x + 3)[x² + 16] = 0

                   (x + 3)(x² + 16) = 0  

Multiplying those using FOIL:

                x³ + 16x + 3x² + 48 = 0

Arranging in descending order:

                x³ + 3x² + 16x + 48 = 0  

Edwin